OpenStudy (anonymous):

Which quadratic equation has its vertex at (-3,5) and opens down? F: y=(x-3)^2+5 H: y=(x+3)^2-5 G: y=-(x+3)^2+5 J: y=-(x-3)^2+5

OpenStudy (rane):

since x = -3 & y = 5 the equation is y = ax^2 + c = -3^2 + 5

OpenStudy (rane):

so your ans is F

OpenStudy (anonymous):

thanks

OpenStudy (rane):

any time :)